OpenStudy (anonymous):

how do you change 2/3+4/5 into a proper fraction or make the denominators the same

OpenStudy (anonymous):

2/3=0.6 , 4/5=0.8 so 0.6+0.8=1.4

OpenStudy (anonymous):

this is grd 7 math

OpenStudy (anonymous):

or : 10/15 + 12/15 = 22/15 = 1.4

OpenStudy (anonymous):

my other examples are 7 1/2 or 19/7 etc

OpenStudy (anonymous):

7 1/2= 15/2

OpenStudy (anonymous):

19/7= 2 5/7

OpenStudy (anonymous):

ik but the question is 2/3 +4/5 add it up for same dinomanator

OpenStudy (anonymous):

2/3 + 4/5 = ( (2*5) + (3*4) ) / (3*5) = 22/15